# FedFDP: Federated Learning with Fairness and Differential Privacy

### v1

@@ Abstract
To address the first issue mentioned above, we propose a federated algorithm with fairness, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, Extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. The code is accessible at https://anonymous.4open.science/r/FedFDP-E754.
1 Introduction
@@
Federated Learning (FL) McMahan et al. (2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Because its ability to overcome data silos by jointly improving model performance without the need to upload local data, FL has gained significant attention in recent year Zhu et al. (2021)Lee et al. (2022). In particular, many studies have made great breakthroughs in model performance Wang et al. (2020); Li et al. (2020a, 2021b) and communication costs Wang et al. (2019); Ling et al. (2023) in FL.
However, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data Zhu et al. (2019) or the membership inference attack Song et al. (2017). Moreover, Melis et al. Melis et al. (2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in the FL, numerous scholars have begun integrating techniques such as differential privacy (DP) Dwork et al. (2014), homomorphic encryption Paillier (1999), and secure multi-party computation Yao (1982) into the federated learning framework. Among these, DP has emerged as the most widely acknowledged approach for privacy protection in federated learning, owing to its minimal computational overhead and rigorous mathematical framework. Many works Abadi et al. (2016); Carlini et al. (2019); Feldman (2020) have demonstrated that by adding proper DP noise in the training phase, the resulted machine learning models can prevent unintentional leakage of private training data, such as membership inference attacks.
@@
Figure 1: The test accuracy of 50 clients utilizing a globally trained model in FL.
Another easily overlooked critical issue in federated learning is fairness. The globally trained model collaboratively by clients may exhibit significant performance disparities across datasets held by different clients. As shown in Fig.1, where the red dots indicate the test accuracy of the global model across various clients after the completion of federated learning using the FedAvg McMahan et al. (2017) algorithm. And we can see the globally trained model applied to individual clients, the performance significant performance disparities across differential clients. More and more scholars have begun to notice this unfair phenomenon and initiated research on it  Wang et al. (2020); Li et al. (2020b).
@@
While numerous articles separately discuss fairness and differential privacy protection in federated learning, there is a lack of research integrating both aspects into federated learning. Balancing model performance, fairness, and differential privacy protection in federated learning presents a challenging problem. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a federated learning framework that ensures both fairness and differential privacy protection, effectively guaranteeing model performance.
Our main contributions are listed below:
  1. 1.
@@
    Firstly, we propose a fairness-aware federated algorithm, denoted as algorithm FedFair by designing a new local loss function that is constructed to simultaneously optimize the two objectives of fairness and model performance (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Fig.1.
  2. 2.
@@
    Building upon FedFair, we propose algorithm FedFDP to achieve differential privacy protection (Section 5) . Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. 3.
@@
    Furthermore, we demonstrate there exist an optimal fairness parameter $\lambda$ value to balance the model performance and fairness through the convergence analysis (Section 6).
  4. 4.
@@
    And then, we perform privacy loss analysis of our algorithm FedFDP by the notion of Rényi differential privacy (RDP) to prove that it satisfies the differential privacy (Section 7).
  5. 5.
@@
    Lastly, through a comprehensive empirical evaluation on three public datasets, we confirm that FedFair and FedFDP outperforms the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2 Related Work
2.1 Fairness in Federated Learning
In the field of machine learning, a multitude of fairness concepts have been advanced (Mehrabi et al. Mehrabi et al. (2021), Yuxin et al. Shi et al. (2021)), each highlighting specific facets and addressing the concerns of diverse stakeholder groups. For instance, Mohri et al. Mohri et al. (2019) advocate for Good-Intent Fairness, while Zhou et al. Zhou et al. (2021) embrace Selection Fairness. Cong and Lyu et al. Cong et al. (2020)Lyu et al. (2020) employ principles from game theory Rabin (1993) to steer the allocation of rewards.
Currently, the predominant fairness criterion in Federated Learning is “Performance Parity”, as proposed by Li et al. Li et al. (2020b). This criterion aims to ensure fairness by evaluating the consistency of performance across different client devices in federated learning. This definition is widely adopted. Zhao et al. Zhao and Joshi (2022) suggested replacing the loss amplification mechanism in Li et al. (2020b) with a straightforward weight reallocation mechanism. Li et al. Li et al. (2021b) focus on optimizing the fairness of local models in personalized federated learning.
2.2 Federated Learning with Differential Privacy
Federated learning with differential privacy has become a common privacy protection method in federated learning. Enayat et al. Ullah et al. (2023) dynamically adjust the compression rate to minimize communication rounds in federated learning by responding to changes in training loss. Zhou et al. Zhou et al. (2023) investigated the effects of local iterations when utilizing the Laplace mechanism for perturbation. Fu et al. Fu et al. (2022) introduced adaptive noise adding, customizing it for different clients and communication rounds based on observed fluctuations in loss values. Ling et al. Ling et al. (2023) determined the optimal number of local epochs through convergence analysis, striking a balance between communication resources and privacy budget. Raouf et al. Kerkouche et al. (2021) improved model accuracy by updating only a fixed subset of model weights while keeping others unchanged.
The aforementioned works each independently address either fairness or privacy protection in federated learning. To the best of our knowledge, we are the first to concurrently consider both fairness and privacy in the context of federated learning.
3 Preliminaries
3.1 Federated Learning
@@
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of Federated Learning is to minimize $F(\mathbf{w})$, in other words, to seek:
$\mathbf{w}^{*}\triangleq\arg\min F(\mathbf{w}).$ (1)
The function $F(\cdot)$ is defined as follows:
$F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$ (2)
@@
3.2 DPSGD
In gradient-based methods, DPSGDAbadi et al. (2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
@@ 3.2 DPSGD
    $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right).$ (4)
  • Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients:
    $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$ (5)
The final step is to proceed with gradient descent as usual. Here, $C$ represents the clipping bound, and $\sigma$ is the noise multiplier. If each client in FL uses DPSGD as the gradient descent method to protect their private data, FL will transform into Differentially Private Federated Learning (DPFL).
3.3 Fairness in Federated Learning
Definition 1.
(Fairness of performance distributionLi et al. (2020b)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally say that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
In this study, we employ the loss of model denote the performance. Equation (6) represents the specific form of fairness in federated learning in this paper, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients.
$\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$ (6)
@@ 19                   $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ and compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
@@
4 Proposed Algorithm: FedFair
In the previous section, we identified the performance objective function as shown in Equation (2) and the fairness objective function as shown in Equation (6). Both of these objective functions are composed of individual client objectives. In this section, we will focus on the performance and fairness of each client and combine these two objectives to construct a new objective function, based on which we propose a fair federated learning algorithm.
We simultaneously optimize performance and fairness, and can get the optimization objective for client $i$ as:
$\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}_{t})\right)^{2},$ (7)
where $F(\mathbf{w}_{t})=\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t-1})$ [^1] The reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training.and $\lambda$ is a hyperparameter used to adjust the degree of fairness, $\lambda\geq 0$. When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness.
The gradient of Equation (7) as follows:
$\nabla H_{i}(\mathbf{w}_{t}^{i})$ $=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (8)
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t})$.
The stochastic gradient descent for Equation (7) as show:
$\mathbf{w}_{t+1}^{i}$ $=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})$
$=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$ (9)

### v2

# FedFDP: Fairness-Aware Federated Learning with Differential Privacy

@@ Abstract.
To address the first issue mentioned above, we propose a fairness-aware federated learning algorithm, termed FedFair. Building upon FedFair, we introduce privacy protection to form the FedFDP algorithm to address the second issue mentioned above. In FedFDP, we devise a fairness-aware clipping strategy to achieve differential privacy while adjusting fairness. Additionally, for the extra uploaded loss values, we present an adaptive clipping approach to maximize utility. Furthermore, we theoretically prove that our algorithm converges and ensures differential privacy. Lastly, extensive experimental results demonstrate that FedFair and FedFDP significantly outperforms state-of-the-art solutions in terms of model performance and fairness. Code and data is accessible at https://anonymous.4open.science/r/FedFDP-5607.
Federated Learning, Differential Privacy, Fairness
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Federated Learning (FL) (McMahan et al., 2017) is a distributed machine learning framework that allows clients to collaboratively train a shared model without revealing their individual datasets. After training their local models, each client transmits only the model parameters, not the raw training data, to a central server. The server aggregates these parameters to update the global model, which is then sent back to clients for further training. Thanks to its ability to overcome data silos by jointly improving model performance without the need to upload raw data, FL has gained significant attention in recent years (Zhu et al., 2021)(Lee et al., 2022). In particular, many studies have made great breakthroughs in model performance (Wang et al., 2020; Li et al., 2020b, 2021b) and communication costs (Wang et al., 2019; Ling et al., 2023) in FL.
@@
Figure 1. The test accuracies of 50 clients utilizing a globally trained model in FL.
[…]
To encourage more clients to participate in federated learning, establishing a fairness-aware federated learning algorithm is necessary (Li et al., 2021c). Current machine learning algorithms that protect fairness, such as  (Dwork et al., 2012; Cotter et al., 2019), have not adequately addressed the issue of performance disparities among clients in federated learning. As shown in Figure 1, the red dots represent the test accuracies of the global model on different clients after completing federated learning with the FedAvg algorithm (McMahan et al., 2017). The global model trained collaboratively by clients exhibits significant unbalanced performance across different datasets held by the various clients. An increasing number of scholars are beginning to pay attention to this performance disparities phenomenon and are starting to work on it  (Li et al., 2020c, 2021b; Zhao and Joshi, 2022).
On the other hand, recent research indicates that trained parameters can leak data privacy, such as the recovery of original data (Zhu et al., 2019) or the membership inference attack (Song et al., 2017). Moreover, Melis et al. (Melis et al., 2022) demonstrated that clients’ training data can be leaked through shared models. In order to enhance data privacy in FL, numerous scholars have begun integrating techniques such as differential privacy (DP) (Dwork et al., 2014), homomorphic encryption (Paillier, 1999), and secure multi-party computation (Yao, 1982) into the federated learning framework. Among these, DP stands out as the premier approach to privacy preservation in federated learning, attributed to its modest computational overhead and robust mathematical underpinnings. This is particularly relevant in defending membership inference attacks, as several studies (Abadi et al., 2016; Carlini et al., 2019; Feldman, 2020) have shown that incorporating appropriate DP during the training phase can effectively prevent inadvertent disclosure of sensitive training information.
@@
Although numerous articles have addressed fairness and differential privacy individually within the context of federated learning, there is a scarcity of research that integrates these two aspects. Achieving a balance between model performance, fairness, and differential privacy protection within federated learning frameworks poses a significant challenge. This paper aims to investigate fairness and differential privacy protection mechanisms in federated learning. By analyzing the relationship between them, we propose a fairness-aware federated learning with differential privacy, effectively guaranteeing model performance when achieving fairness and privacy protection.
Our main contributions are listed as follows:
  1. (1)
@@
    Firstly, a fairness-aware federated algorithm is proposed, denoted as FedFair. It incorporates a novel loss function, designed to optimize the objectives of fairness and model performance simultaneously (Section 4). FedFair algorithm achieves the excellent fairness, as shown by the green dots in Figure1.
  2. (2)
@@
    Building upon FedFair, an algorithm referred to as FedFDP is proposed to achieve differential privacy protection (Section 5). Particularly, we device a fairness-aware gradient clipping strategy to ensures both differential privacy and allows for the adjustment of fairness in federated learning (Section 5.1). Additionally, we propose an adaptive clipping approach method for the additional loss values uploaded by each client to achieve optimal utility (Section 5.2).
  3. (3)
@@
    Furthermore, it is demonstrated that an optimal fairness parameter $\lambda$ is found, which, when determined, enables the fastest convergence of FedFDP in balancing model performance and fairness through the convergence analysis (Section 6).
  4. (4)
@@
    And then, a privacy loss analysis is conducted on algorithm FedFDP using the notion of Rényi differential privacy (RDP) to verify its satisfaction of differential privacy requirements (Section 7).
  5. (5)
@@
    Lastly, through a comprehensive empirical evaluation on three public datasets, it is confirmed that FedFair and FedFDP outperform the state-of-the-art solutions in terms of model performance and fairness (Section 8).
2. Related Work
Differentially private federated learning (DPFL) is widely used for privacy protection in federated learning. Many studies have focused on the trade-off between model performance, communication efficiency, and privacy budget in DPFL (Ullah et al., 2023; Zhou et al., 2023; Fu et al., 2022; Ling et al., 2023; Kerkouche et al., 2021; Fu et al., 2024). In addition that, various fairness concepts in machine learning have been proposed, each emphasizing different aspects and addressing the needs of different stakeholders. For example, concepts like Good-Intent Fairness (Mohri et al., 2019) and Selection Fairness (Zhou et al., 2021) have been introduced. Further discussions on these concepts can be found in additional literature (Cong et al., 2020; Lyu et al., 2020; Mehrabi et al., 2021; Shi et al., 2021).
At present, the primary criterion for fairness in FL is group fairness, which aims to mitigate biases associated with sensitive attributes like gender and race. Relevant research has shown that in federated learning with differential privacy, adjusting parameters such as the clipping threshold and noise multiplier can help enhance the algorithm’s group fairness (Gu et al., 2022). There is a growing body of literature that addresses the concept of group fairness specifically within the framework of DPFL (Gálvez et al., 2021; Padala et al., 2021; Gu et al., 2022, 2020; Pentyala et al., 2022).
The more predominant fairness criterion in FL is balanced performance fairness, as introduced by Li et al. (Li et al., 2020c). This approach ensures fairness by assessing the uniformity of model performance across participating clients. The concept has gained broad acceptance. Building on this, Zhao et al. (Zhao and Joshi, 2022) proposed an alternative to the loss amplification mechanism, advocating for a direct weight reallocation strategy. Concurrently, Li et al. (Li et al., 2021b) concentrated on enhancing the fairness of local models within personalized FL. Our study pioneers the investigation of balanced performance fairness within the context of DPFL.
In comparison to definitions such as group fairness, balanced performance fairness is more meaningful for data providers such as enterprises and institutions, moreover, it can effectively motivate clients participate in the federated system. However, achieving balanced performance fairness typically requires adjusting the direction of model updates. It is a challenge to refine the model for balanced performance fairness while integrating DP techniques, such as clipping and noise addition, into the model’s update process. To the best of our knowledge, we are the first to consider the work of balanced performance fairness in DPFL.
3. Preliminaries
Table 1. Summary of main notations
$F(\mathbf{w}$) Global loss function
$F_{i}(\mathbf{w}$) Local loss function for client $i$
$T$ Communication rounds
$t$ Communication round index
$D_{i}$ Dataset of client $i$
$D$ The union of $D_{i}$
$\mathbf{w}_{t}^{i}$ Local model parameter at node $i$ at round $t$
$\mathbf{w}^{*}$ Optimal model parameter that minimizes $F(\mathbf{\mathbf{w}})$
$\eta$ Gradient descent step size
$\lambda$ Fairness parameter
$N$ Total number of clients
$q$ batch sample ratio
$|\mathcal{B}_{i}|$ Batch size of client $i$, with the mathematical expectation equal to $|D_{i}|\cdot q$
$p_{i}$ Weight of client $i$, equals to $|D_{i}|/|D|$
$\xi_{j}$ Single data sample in $D_{i}$, indexed by $j$
$\sigma$ Noise multiplier for gradient
$\sigma_{l}$ Noise multiplier for loss
$C$ Original clipping bound for gradient
$C_{l}$ Original Clipping bound for loss
|||||$\cdot$||| | ⋅ | | $L_{2}$-norm
3.1. Federated Learning and Fairness
@@
We use $\mathbf{w}^{i}_{t}$ signifies the model acquired from client $i$ at iteration $t$, while $\mathbf{w}_{t}$ denotes the server-aggregated model. The objective of FL is to minimize $F(\mathbf{\cdot})$, in other words, to seek:
(1) $\mathbf{w}^{*}$ $\triangleq\arg\min F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}).$
@@
We have defined balanced performance fairness in FL as follows. Specifically, this type of fairness is defined based on the performance variance of the model across clients: the smaller the variance, the fairer the model.
Definition 0.
(Balanced performance fairness (Li et al., 2020c)) For trained models $\mathbf{w}_{1}$ and $\mathbf{w}_{2}$, we informally recognize that model $\mathbf{w}_{1}$ provides a more fair solution to the federated learning objective in Equation (1) than model $\mathbf{w}_{2}$ if the performance of model $\mathbf{w}_{1}$ on the $N$ devices is more uniform than the performance of model $\mathbf{w}_{2}$ on the $N$ devices.
3.2. Differential Privacy
Differential privacy is a formal mathematical model that quantifies privacy leakage in data analysis algorithms. It stipulates that alterations to a single record within the training data should not induce significant statistical variations in the algorithm’s results.
Definition 0.
(Differential Privacy (Dwork et al., 2014)). $(\epsilon,\delta)$-differential privacy is achieved by a randomized mechanism $\mathcal{M}:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, for any two neighboring databases $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ that differ in only a single entry, and $\forall$ $\mathcal{S}\subseteq\text{Range}(\mathcal{R})$:
(2) $\operatorname{Pr}\left[\mathcal{M}\left(D_{i}\right)\in\mathcal{S}\right]\leq e^{\epsilon}\operatorname{Pr}\left[\mathcal{M}\left(D_{i}^{\prime}\right)\in\mathcal{S}\right]+\delta.$
By adding random noise, we can achieve differential privacy for a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$ according to Definition 2. The $l_{k}$-sensitivity determines how much noise is needed and is defined as follow. We use $l_{2}$ sensitivity as in the previous work.
Definition 0.
($\mathbf{l_{k}}$-Sensitivity(Dwork et al., 2006)) For a function $f:\mathcal{X}^{n}\rightarrow\mathbb{R}^{d}$, we define its $l_{k}$ norm sensitivity (denoted as $\Delta_{k}f$) over all neighboring datasets $D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}$ differing in a single sample as
(3) $\text{sup}_{D_{i},D_{i}^{\prime}\in\mathcal{X}^{n}}||f(D_{i})-f(D_{i}^{\prime})||_{k}\leq\Delta_{k}f.$
DPSGD. In gradient-based methods, DPSGD(Abadi et al., 2016) is widely used in privacy preservation. It performs per-sample gradient clipping and noise addition:
  1. (1)
@@ 3.2. Differential Privacy
    (5) $\hat{g}^{i,j}_{t}=g^{i,j}_{t}/\max\left(1,\frac{\|g^{i,j}_{t}\|}{C}\right),$
    where $C$ represents the clipping bound.
  3. (3)
    Thirdly, adding Gaussian noise with mean 0 and standard deviation $C\sigma$ to the sum of clipped gradients, where $\sigma$ is the noise multiplier:
    (6) $\tilde{g}^{i}_{t}=\frac{1}{|\mathcal{B}_{i}|}\left(\sum_{j=1}^{|\mathcal{B}_{i}|}\hat{g}^{i,j}_{t}+C\sigma\cdot\mathcal{N}(0,\mathbf{I})\right).$
The final step is to proceed with gradient descent as usual. When each client in FL utilizes DPSGD as the optimization algorithm to safeguard their private data, the framework transitions from FL to DPFL.
@@ 19             $\mathbf{w}_{t}^{i}=\mathbf{w}_{t}^{i}-\eta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i},b)$;
21      $\mathbf{w}_{t+1}^{i}\leftarrow\mathbf{w}_{t}^{i}$ , compute $F_{i}(\mathbf{w}_{t+1}^{i})$;
@@
4. Proposed Algorithm: FedFair
We previously defined the performance objective in Equation (1) and balanced performance fairness in Definition 1. Here, we use these to develop a new objective function that informs our proposed fairness-aware federated learning algorithm.
In this study, we employ the loss of model denote the performance. Equation (7) represents the specific form of fairness in federated learning, which is the weighted variance of the loss values of global trained model $\mathbf{w}$ across all clients. Essentially, the smaller the value of $\Psi$, the fairer the $\mathbf{w}$ for all clients:
(7) $\Psi(\mathbf{w})\triangleq\sum_{i=1}^{N}p_{i}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
Upon incorporating Equation (7) into Equation (1), we introduce a comprehensive objective function that integrates considerations of fairness:
(8) $\min_{\mathbf{w}}H(\mathbf{w})=F(\mathbf{w})+\frac{\lambda}{2}\sum_{i=1}^{N}\left(F_{i}(\mathbf{w})-F(\mathbf{w})\right)^{2}.$
For client $i$ during round $t$, the refined objective is articulated as follows:
(9) $\min_{\mathbf{w}^{i}_{t}}H_{i}(\mathbf{w}^{i}_{t})=F_{i}(\mathbf{w}^{i}_{t})+\frac{\lambda}{2}\left(F_{i}(\mathbf{w}^{i}_{t})-F(\mathbf{w}^{i}_{t})\right)^{2},$
where $F(\mathbf{w}^{i}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}^{i}_{t})$ and $F_{i}(\mathbf{w}^{i}_{t})$ is determined using the training dataset [^1] We use $F(\mathbf{w}_{i}^{t})\leftarrow F(\mathbf{w}_{t})\triangleq\sum_{i=1}^{N}p_{i}F_{i}(\mathbf{w}_{t}^{i})$, the reason for $F(\mathbf{w}_{t})\neq F(\mathbf{w}_{t},\xi_{test})$ is that the test data $\xi_{test}$ from the server cannot be used for training. from round $t-1$, as delineated in line 1 of Algorithm 1, and the $\lambda$ ($\lambda\geq 0$) is a hyperparameter used to adjust the degree of fairness  [^2] When $\lambda=0$, the algorithm becomes traditional federated learning. When $\lambda\rightarrow\infty$, the algorithm only focus on fairness. However, if $\lambda$ is too large, it will cause very steep gradients which might prevent the model from converging. Fortunately, we theoretically found an optimal value for $\lambda$ in section 6.2. In addition, as shown in Figure 3, we experimentally confirmed the existence of the optimal $\lambda$.. Consequently, within the context of Equation (9), $F(\mathbf{w}^{i}_{t})$ is considered a constant when the gradient is being computed. So, the gradient of Equation (9) as follows:
(10) $\nabla H_{i}(\mathbf{w}_{t}^{i})=\nabla F_{i}(\mathbf{w}_{t}^{i})+\lambda\Delta_{i}\cdot\nabla F_{i}(\mathbf{w}_{t}^{i})=(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i}),$
where $\Delta_{i}=F_{i}(\mathbf{w}_{t}^{i})-F(\mathbf{w}_{t}^{i})$.
The stochastic gradient descent by using Equation (10) as show:
(11) $\mathbf{w}_{t+1}^{i}=\mathbf{w}_{t}^{i}-\eta(1+\lambda\Delta_{i})\nabla F_{i}(\mathbf{w}_{t}^{i})=\mathbf{w}_{t}^{i}-\eta_{i}\nabla F_{i}(\mathbf{w}_{t}^{i}),$